OpenStudy (anonymous):

equation for a hyperbola with vertices: (16,0) and (-16,0), conjugate axis of length 10 units

OpenStudy (anonymous):

a = 16; b= 10/2 = 5 . So the hyperbola will be (x^2 / 16^2 ) - (y^2/5^2) = 1